OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−4, −1) and (4, 5)

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] find m which is the slope m=3/4 \[y+1=\left( 3/4 \right)(x+4)\] now solve for y y=(3/4)x+2 so the coordinate for the y intercept is (0,2)